Activity 2. Construct Me!




A population consists of the five numbers 2, 3, 6, 8 and 11.




Consider samples of size 2 that can be drawn from this population.




a. List all the possible samples and the corresponding mean.




Sample




Mean




19




b. Construct the sampling distribution of the sample means.




Sample Mean




Probability




Frequency








c. Draw a histogram of the sampling distribution of the




meets.

1
Expert's answer
2022-05-06T06:00:06-0400

a. We have population values 2, 3, 6, 8 and 11 population size N=5 and sample size n=2.

Mean of population "(\mu)" = 

"\dfrac{2+3+6+8+11}{5}=6"


Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}"


"=\dfrac{16+9+0+4+25}{5}=10.8"

"\sigma=\sqrt{\sigma^2}=\sqrt{10.8}\approx3.2863"

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{5}C_2=10."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 2,3 & 5/2 \\
 \hdashline
 2 & 2,6 & 8/2 \\
 \hdashline
 3 & 2,8 & 10/2 \\
 \hdashline
 4 & 2,11 & 13/2 \\
 \hdashline
 5 & 3,6 & 9/2 \\
 \hdashline
 6 & 3,8 & 11/2 \\
 \hdashline
 7 & 3,11 & 14/2 \\
 \hdashline
 8 & 6,8 & 14/2 \\
 \hdashline
 9 & 6,11 & 17/2 \\
 \hdashline
 10 & 8,11 & 19/2 \\
 \hdashline
\end{array}"



b.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) &\bar{X}^2 f(\bar{X})\\ \hline
 5/2 & 1/10 & 5/20 & 25/40\\
 \hdashline
 8/2 & 1/10 & 8/20 & 64/40\\
 \hdashline
 9/2 & 1/10 & 9/20 & 81/40\\
 \hdashline
 10/2 & 1/10 & 10/20 & 100/40\\
 \hdashline
 11/2 & 1/10 & 11/20 & 121/40\\
 \hdashline
 13/2 & 1/10 & 13/20 & 169/40\\
 \hdashline
 14/2 & 2/10 & 28/20 & 392/40\\
 \hdashline
 17/2 & 1/10 & 17/20 & 289/40\\
 \hdashline
 19/2 & 1/10 & 19/20 & 361/40\\
 \hdashline
\end{array}"


Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=6=\mu"


The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{1602}{40}-(6)^2=\dfrac{162}{40}=\dfrac{81}{20}= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

"\sigma_{\bar{X}}=\sqrt{\dfrac{81}{20}}\approx2.01246"

c.